### v1

# Look into the Future: Deep Contextualized Sequential Recommendation

Abstract.
Sequential recommendation focuses on mining useful patterns from the user behavior history to better estimate his preference on the candidate items. Previous solutions adopt recurrent networks or retrieval methods to obtain the user’s profile representation so as to perform the preference estimation. In this paper, we propose a novel framework of sequential recommendation called Look into the Future (LIFT), which builds and leverages the contexts of sequential recommendation. The context in LIFT refers to a user’s current profile that can be represented based on both past and future behaviors. As such, the learned context will be more effective in predicting the user’s behaviors in sequential recommendation. Apparently, it is impossible to use real future information to predict the current behavior, we thus propose a novel retrieval-based framework to use the most similar interaction’s future information as the future context of the target interaction without data leakage. Furthermore, in order to exploit the intrinsic information embedded within the context itself, we introduce an innovative pretraining methodology incorporating behavior masking. This approach is designed to facilitate the efficient acquisition of context representations. We demonstrate that finding relevant contexts from the global user pool via retrieval methods will greatly improve preference estimation performance. In our extensive experiments over real-world datasets, LIFT demonstrates significant performance improvement on click-through rate prediction tasks in sequential recommendation over strong baselines. The experiment code is provided at https://anonymous.4open.science/r/LIFT-277C/.
Sequential Recommendation, Context Representation, Retrieval-Enhanced Methods, Pretraining
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06conference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management ; October 21 –– 25 2024 ; Barcelona, Spain
1. Introduction
Deep learning has been widely adopted for predicting user behaviors in online recommender systems (Zhang et al., 2021). For the scenarios with sequences of user behaviors, sequential recommendation techniques, which extract valuable information from the past behavior sequence of the target user to predict his next behavior, have been well studied (Pi et al., 2019; Huang et al., 2018).
As illustrated in Figure 1, the major techniques for deep learning-based sequential recommendation are generally threefold. Firstly, deep architectures are designed to extend the capability of traditional collaborative filtering models for better mining of the feature interactions (Guo et al., 2017; Zhang et al., 2016; Wang et al., 2017). Secondly, different models like recurrent neural networks (RNN), memory networks, or transformers are adopted to learn an effective representation of the user behavior sequence, based on which a deep learning predictor is built to predict the behavior label for the current-step recommendation (Pi et al., 2019; Chen et al., 2018; Sun et al., 2019). Dealing with the long behavior sequence problem, thirdly, the retrieval methods are leveraged to fetch far early but relevant behaviors, which are then aggregated according to the target prediction condition and fed into the final label predictor (Qi et al., 2020; Qin et al., 2020).
However, the retrieval-enhanced models for sequential recommendation still retrieve in the “raw” data space, i.e., the individual behavior data, which actually loses the contextual information of the user making that particular behavior. Users’ behaviors over the recommended items commonly exhibit interdependence. Typically, a user browses items with a specific intent, and within a certain time window, his attention is often drawn to items within a particular theme. For example, when a user is browsing ketchup, they may also browse many other kitchen condiments before and after. To this end, we posit the concept of a user’s browsing sequence encompassing historical and future behaviors as the context of that behavior. The incorporation of contextual information pertaining to the future behavior enhances the effectiveness of modeling user behaviors. Similarly, in the area of natural language processing (NLP), to model the word representation, the words before and after the target word are often used (Mikolov et al., 2013; Devlin et al., 2018). Instead, in the field of recommendation systems, the future behavior sequences are often overlooked (Zhou et al., 2018b, 2019) because the future information is inaccessible when predicting the current behavior. However, we may obtain insight into future user behavior by identifying analogous contexts from the observed data of other users and employing them to better model the context of the target user.
In this paper, we propose a novel framework of sequential recommendation called Look into the Future (LIFT), which focuses on retrieval in the user contextual information space while leveraging the future behaviors of the retrieved context, as illustrated on the bottom part of Figure 1. LIFT works in scenarios where the behavior sequences of multiple users are accessible. In sequential recommendation, the representation of context interactions plays an important role in behavior prediction. All previous works in recommendation focus on modeling the historical part of the context. From a hindsight perspective, the entire context of the target interaction should include future interactions. As mentioned above, it is impossible to use the real future information to predict the current behavior. Thus, given the current-step context with the candidate item for behavior label prediction, LIFT performs retrieval to fetch the most similar interaction behavior from the whole user pool. Then, extending from each of the retrieved behaviors, the historical and future behavior data are imported to enrich the context of that retrieved behavior. Note that each of the included ‘future’ behaviors is still temporally earlier than the current timestep of the target user behavior, which avoids any data leakage issue. Such ‘future’ behaviors can be regarded as a kind of privileged information (Vapnik et al., 2015; Xu et al., 2020), which is not accessible for current-step prediction but is accessible for historical behavior data. To our knowledge, there is no previous work on sequential recommendation that performs effective retrieval of multiple users’ behavior contexts.
Figure 1. The comparison between LIFT and conventional models entails several key distinctions: a) Traditional models rely solely on instant user and item information when making predictions. b) Sequential models, conversely, typically incorporate the user’s historical interactions to capture their evolving interests over time. c) Retrieval-based models perform retrieval to fetch far-before but relevant historical behaviors to build the user profile for predictions. d) LIFT focuses on interaction context, encompassing both the historical and future sequence of interactions for each user-item interaction. LIFT’s primary objective is to exploit this comprehensive context information to enhance prediction accuracy. However, note that in practical scenarios, accessing genuine future information is unattainable. To overcome this limitation, we leverage the future sequences of relevant observed interactions as a surrogate to approximate the future context for the current interaction. Note that the included ‘future’ interactions are still temporally earlier than the current target behavior for prediction, which avoids any data leakage issue.
Furthermore, it is crucial not only to utilize information derived from the target sample but also to exploit the inherent information within the context itself for the purpose of enhancing contextual representation. Therefore, besides the supervised training with the label prediction loss, a representation learning loss is much important in our task. To learn an effective representation of the context with both history and future behavior data, we further devise a pretraining method that performs masked behavior prediction during the pretraining stage. We conduct extensive experiments on three real-world sequential recommendation datasets, where LIFT demonstrates significant performance improvements over strong baselines.
Overall, the main contributions of this paper are threefold.
  • We propose to incorporate the future information as part of the context information. We propose a novel framework that leverages retrieval techniques to use relevant context information to enhance the prediction performance of sequential recommendations while still avoiding data leakage. To our knowledge, this is the first work that makes the most use of the entire context information in the recommendation system based on retrieval techniques.
  • In the pursuit of obtaining valuable representations of retrieved user behaviors, we adopt an approach that leverages intrinsic contextual information. Specifically, we devise a pretraining methodology that is tailored to the format of user sequence data and introduces a novel self-supervised loss function referred to as mask behavior loss.
  • To prevent noise introduced by the retrieved data, we introduce a key-based attention aggregation technique aimed at consolidating contextual representations in accordance with the specific target sample and the keys contained within the retrieved contextual information.
2. Related Work
Sequential Recommendation. For sequential recommendation, user behavior modeling is the core technique that mines user preferences from historical interaction behaviors meticulously. To better extract informative knowledge from user’s behavior sequence, various network structures have been proposed, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Attention Networks, and Memory Networks. GRU4Rec (Hidasi et al., 2015) designs Gated Recurrent Units (GRUs) to capture the preference-evolving relationship while Caser (Tang and Wang, 2018) leverages the horizontal and vertical convolution to model skip behaviors at both union-level and point-level. Moreover, the attention mechanism is the most popular method for modeling item dependencies, and several influential works are proposed, including SASRec (Kang and McAuley, 2018), DIN (Zhou et al., 2018b), DIEN (Zhou et al., 2019), and BERT4Rec (Sun et al., 2019). Among these, DIN and DIEN leverage the target-attention network to identify important historical items, while SASRec and BERT4Rec apply the self-attention network with Transformer architecture to excavate behavior dependencies. Besides, memory-based methods (Wang et al., 2018; Pi et al., 2019) are also proposed to store user behavior representations in a read-write manner. To better mine user preferences from their historical behaviors, two kinds of work are proposed to further extend the behavior modeling. The first one is multi-behavior modeling (Zhou et al., 2018a; Yuan et al., 2022) that explicitly leverages different types of behavior (e.g., click and purchase behaviors) to measure item correlations within different behavior sequences, thus capturing users’ diverse interests. The second one is behavior modeling with side information (Zhang et al., 2019) that involves various item attribute features (e.g., category) in addition to item IDs for better exploiting the rich knowledge.
Retrieval-enhanced Recommendation. To enhance the performance of recommender systems, the retrieval-enhanced recommendation is proposed, where the most relevant items are retrieved from an extremely long behavior sequence. Specifically, UBR4CTR (Qin et al., 2020) and SIM (Qi et al., 2020) are designed to retrieve beneficial behaviors from the user’s historical behavior, where UBR4CTR deploys the search engine method while SIM uses the hard search and soft search approaches. To make the search procedure end-to-end, ETA (Chen et al., 2021) is proposed by leveraging the SimHash algorithm to map the user behavior into a low-dimensional space, hence achieving learnable retrieval. Moreover, recent works further extend the retrieval-enhanced recommendation from item-level retrieval to sample-level retrieval. RIM (Qin et al., 2021) is the first to deploy this method that leverages the search engine to retrieve several relevant samples from the search pool and performs neighbor aggregation.
Pretraining for Recommendation. Pretraining the deep learning models (or the data representation) with self-supervised learning methods has been widely studied in natural language processing and computer vision. Generally, there are two major categories of self-supervised training methods, namely contrastive learning (Gao et al., 2021; Chen et al., 2020) and mask recovery (Brown et al., 2020; Jacob et al., 2019; He et al., 2022). For recommender systems or tabular data prediction tasks, there are some recent attempts in this direction. To list a few examples, BERT4Rec (Sun et al., 2019) focuses on learning the representation of sequential behaviors via masking the final item of a subsequence of user behaviors. SCARF (Bahri et al., 2022) raises a contrastive learning loss via feature corruption over the tabular data. MISS (Guo et al., 2022) proposes interest-level contrastive losses to take the place of sample-level losses in order to mine self-supervision signals from user behaviors of multiple interests. Also, there are recent attempts to leverage the knowledge from the pretraining work on outsourced data, such as knowledge graphs (Wong et al., 2021) and language corpus (Cui et al., 2022), to enhance the recommender systems.
3. Formulation & Preliminaries
Notation Description
$\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{test}}$, $\mathcal{D}_{\text{retrieval}}$ Training set, test set, retrieval set
$x_{z}$, $\mathbf{x}_{z}$ The raw feature and the embedding of $z$-th sample
$x_{t}$, $\mathbf{x}_{t}$ The raw feature and the embedding of target sample
$E_{\omega}$ The encoder and its parameters
$R$ The retriever
$F_{\theta}$ The predictor and its parameters
$d$ An interaction
$v$ The encoder output vector dimension
$w$ The feature embedding vector dimension
$K$ The size of retrieved contexts
$L$ The length of history and future sequence
$c$, $h$, $f$ The context sequence, history sequence, future sequence
$C$, $H$, $F$ The context sequence set, history sequence set, future sequence set
Table 1. Notations and corresponding descriptions.
In this section, we formulate the studied problem and introduce the preliminaries. The notations and their descriptions are summarized in table 1. A recommendation dataset $\mathcal{D}=\{d_{z}\}_{z=1}^{N}$ is represented as a set of interactions $d_{z}$ between user $u_{z}$ and item $i_{z}$. Each interaction $d_{z}$ can be formulated as the feature-label pair $(x_{z},y_{z})$, where $x_{z}=\{x_{j}^{z}\}_{j=1}^{M}$ contains the user and item features and a label $y\in\{0,1\}$ indicating whether the user will click the item or not, i.e., click-through rate (CTR).
A user’s interaction sequence $s$ can be defined as a list of consecutive interactions for the same user sorted by time $s=[d_{1},d_{2},\ldots,d_{T}]$ where $d_{1}$ is the earliest interaction and $d_{T}$ is the latest one. We define the last $L$ interactions before $d_{z}$ as the history sequence $h_{z}$ and the future $L$ interactions after $d_{z}$ as the future sequence $f_{z}$. Then we define the full context $c_{z}$ as the combination of the history interaction sequence $h_{z}$ and future interaction sequence $f_{z}$:
(1) $c_{z}=(h_{z},f_{z})~{}.$
To represent $d_{z}$, previous works on sequential recommendation only consider modeling the historical part of the context (Hidasi et al., 2015; Zhou et al., 2018b), while in this work, we use the full context to model $d_{z}$.
The goal of a traditional recommendation task is to predict the label of the interaction $d_{z}$ based on the features $x_{z}$ from users, items, and the user $u_{z}$’s history behavior sequences. Such a prediction can be formulated as
(2) $\hat{y}_{z}=F_{\theta}(x_{z},h_{z})~{},$
where $F_{\theta}$ is the learning model with the parameter $\theta$.
Here, we would use both the historical part information $h_{z}$ and future part information $f_{z}$. However, the future information for $d_{z}$ is not visible in the inference stage. So we design a retrieval-based framework to retrieve the most relevant interactions’ future sequence as the future sequence for the interaction $d_{z}$. Moreover, we still use the most similar interaction’s history sequence to enhance the historical information.
Following the experiment setting from other retrieval-based sequential recommendation methods (Qin et al., 2021), we split the dataset as $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}},\mathcal{D}_{\text{retrieval}}$ for training, testing, and retrieval, respectively. With the relevant contexts sequences, which include future sequences and history sequences, the label prediction is formulated as
(3) $\hat{y}_{z}=F_{\theta}(x_{z},h_{z},C_{z})~{},$
where $C_{z}$ represents the context sequences for $d_{z}$ and
(4) $C_{z}=R(x_{z})~{},$
where $R$ is a retriever that could fetch relevant context sequences according to $x_{z}$. We first retrieve relevant interactions $\{d_{r}\}$. Then use the retrieved interactions to find their contexts to build $C_{z}$. We will describe the procedure details in the next section.
Traditional methods only use the label from the target interaction $d_{z}$ to train $F_{\theta}$. However, if we only use the labels from the target interaction $d_{z}$, we will neglect most of the behavior information in the retrieved context sequences $C_{z}$. Inspired by pretrained language models (Devlin et al., 2019; Brown et al., 2020), which have achieved tremendous success in the field of natural language processing (NLP) by learning universal representations in a self-supervised manner, we propose a self-supervised pretraining method that is supervised by the signal from the retrieved sequences itself. We denote the pretrained encoder as $E_{\omega}$ where the input of the encoder is any user interaction sequence $s$ and the output of $E_{\omega}$ is a $v$-dimension vector. The embedding of any user interaction sequence $s$ is written as
(5) $\mathbf{s}=E_{\omega}(s)~{},$
where $s\in\mathbb{R}^{1\times v}$ and the retrieved context sequences set includes a history sequence set $H$ and a future sequence set $F$. We retrieve $K$ context sequences, where $K$ can be tuned in this framework. Then the retrieved context embedding set is written as
(6) $\mathbf{C}=(\mathbf{H},\mathbf{F})=(E_{\omega}(H),E_{\omega}(F))~{},$
where $H\in\mathbb{R}^{K\times v},F\in\mathbb{R}^{K\times v},C\in\mathbb{R}^{K\times 2v}$.
As such, the parameters can be optimized by minimizing the self-supervised loss function as
[…]
where $\mathcal{S}_{\text{pretrain}}$ is a sequence set sampled from the retrieved dataset $\mathcal{D}_{\text{retrieval}}$. As the retriever $R$ is non-parametric, we optimize $F_{\theta}$ and $E_{\omega}$ independently. Specifically, after pretraining and fixing the parameter of $E_{\omega}$, we train $F_{\theta}$ via
[…]
where $\mathbf{h}_{z}$ is the history sequence of $u_{z}$, encoded by $E_{\omega}$.
Figure 2. The architectural components of the encoder and predictor within the LIFT framework are as follows: (a) Pretrained Sequence Encoder: The embedding layer is omitted in this component. LIFT employs a decoder-only Transformer architecture as the encoder, which undergoes pretraining via the mask behavior loss. During the pretraining stage, the primary focus is on leveraging contextual information inherent within the sequence data itself. (b) Training of the Predictor: During the training stage, emphasis is placed solely on the training of the predictor. LIFT incorporates three distinct types of information to inform its predictions, namely the target sample $x_{t}$ itself, the user’s historical interactions, and the retrieved context, encompassing historical interactions from similar instances as well as future interactions. In this phase, the label information utilized for training is exclusively derived from the target samples.
4. The LIFT Framework
In this section, we provide an overview of the whole LIFT framework and then describe the details of its the specific descriptions for its component encoder, retriever, and predictor, respectively.
Figure 3. The overview of major components and workflow in LIFT. In the initial phase (Stage 1), we extract a sequence dataset from the original dataset, subsequently employing this sequence dataset for the purpose of pretraining the encoder. In the subsequent phase (Stage 2), we establish a datastore utilizing the dataset $\mathcal{D}_{\text{retrieval}}$. Following this, we proceed to train the predictor denoted as ”$Predictor$” using the dataset $\mathcal{D}_{\text{train}}$.
4.1. Overview
There are three major components in LIFT, i.e., the encoder, retriever, and predictor. The encoder $E$ is responsible for encoding the user sequence to a $v$-dimensional vector. Given target interaction $d_{t}$, its feature part $x_{t}$ is regarded as a query, and the label part $y_{t}$ is to be predicted. The retriever $R$ is used to retrieve relevant context sequences $C_{t}$. The predictor $F_{\theta}$ makes use of the feature vector $x_{t}$, the user $u_{t}$’s history sequence $h_{t}$, and the relevant context sequence set $C_{t}$ to predict the final label $\hat{y_{t}}$.
There are two steps in LIFT framework: pretraining and predicting. During pretraining, an encoder $E$ is trained on a sampled $\mathcal{S}_{\text{pretrain}}$ from $\mathcal{D}_{\text{retrieval}}$ For $predicting$, the LIFT model is first using the retriever $R$ to find the top-$K$ relevant user context sequences $C_{t}$. And then, each sequence in $C_{t}$ will be encoded by $E_{\omega}$. After aggregating the information of the target user and item $x_{t}$, the target user’s history $h_{t}$, retrieving sequence $C_{t}$, the predictor $F_{\theta}$ will generate the output label prediction.
The main contribution of the LIFT framework is that it makes use of the full user context to help predict the target user behavior. Another distinctive feature of LIFT is that it leverages the pretraining method in a quite new paradigm and significantly improves the prediction performance in our experiment, which may open a new research direction of the sequential recommendation task. Next, we will describe the three components of the LIFT framework in detail.
4.2. Encoder
The encoder $E_{\omega}$ maps a user interaction sequence into a real-valued vector. It is pretrained separately in the framework.
Model Architecture. LIFT adopts a decoder-only Transformer as the encoder, which is based on the original implementation described in Vaswani et al. (2017). Transformer has been shown to perform strongly in language modeling (Brown et al., 2020; Devlin et al., 2019), machine translation (Vaswani et al., 2017), recommender system (Sun et al., 2019), etc. Different tasks may leverage different parts of the Transformer architecture. The difference between the Transformer encoder and decoder is that the self-attention sub-layer in the decoder is modified to make sure that the prediction in position $p$ could only depend on the positions earlier than $p$ (Vaswani et al., 2017). For the data of the recommender system exhibiting temporal characteristics, we use the decoder part of the Transformer. We also compare the performance of different encoders in the experiment, which shows the superiority of the Transformer decoder.
Input/Output. The input of the encoder $E_{\omega}$ is a user interaction sequence $s=\{d_{1},d_{2},\ldots,d_{T}\}$. The output of $E_{\omega}$ is an embedding representing the sequence.
To feed the interaction $d_{z}$ into the encoder, $d_{z}$’s feature part $x_{z}$ is fed into an embedding layer. The embedding layer will map the one-hot encoded vector $x_{z}$ into a dense vector $\bf{x}_{z}$, where each feature $a_{i}$ in $x_{z}$ will be mapped into a $w$-dimensional vector and then be concatenated into one $(M\times w)$-dimensional dense vector. The label part of the $d_{z}$ will be preserved. In the training stage, if the interaction $d_{z}$ is chosen to be masked, the label will be replaced by a special token [MSK]. We will describe the training loss in the next subsection. After processing the sequence, the output will be sent into the decoder-only Transformer decoder sequentially, and the output is a set of hidden states $O=\{o_{1},o_{2},...,o_{T}\}$. We choose the last state $o_{T}$ as the representation of the interaction sequence $\bf{s}$.
Data. The pretraining dataset $\mathcal{S}_{\text{pretrain}}$ is sampled from $\mathcal{D}_{\text{retrieval}}$. For each user $u$ in $\mathcal{U}$, we use all the interactions in $\mathcal{D}_{retrievel}$ of $u$ to assemble an interaction sequence $s_{u}$. Then we get a sequence set $S$. For any sequence $s$ in $S$, we pick up a subsequence every $L$ interactions. Using these subsequences, we get the pretraining dataset $\mathcal{S}_{\text{pretrain}}$.
Mask Behavior Loss. To leverage the inherent information contained within the context, we formulate a self-supervised loss function. In contrast to certain existing sequential recommendation models (Zhou et al., 2018b, 2019), wherein only the target interaction label is considered as the prediction target for any given interaction sequence, all other interaction labels within the sequence are omitted. It is rational to argue that incorporating more behavioral information from the context as the supervision signal would lead to a more comprehensive representation of the context. Consequently, such an approach is expected to yield greater benefits for the prediction of target behavior.
As shown in Figure 2, the left part describes the encoder architecture and loss in the pretraining procedure. In order to train a deep representation of the sequence, we mask part of the interaction’s labels and then predict those masked behavior. Unlike the Cloze task (Taylor, 1953) in natural language processing, it is difficult to predict the masked tokens in the user interaction sequences because there is no syntax in the recommendation. An interaction $d_{i}$ in the input sequence $s$ can be represent as
(9) $d_{i}=\left\{\begin{matrix}(x_{i},\texttt{MSK})&\text{if $d_{i}$ is a masked behavior}\\ (x_{i},y_{i})&\text{otherwise}\end{matrix}\right.$
where $i\in[1,T]$.
We adopt the binary cross entropy loss as the objective function. In the pretraining procedure, we feed the masked interactions’ corresponding output as input into a multi-layer perceptron (MLP) to predict the masked label. Thus the pretrain loss $\mathcal{L}_{\text{pretrain}}$ can be represented as $\mathcal{L}_{\text{pretrain}}=$
(10) $-\sum_{d_{i}\in\mathcal{M}}\left[y_{i}\log\sigma(\text{MLP}(o_{i}))+(1-y_{i})\log(1-\sigma(\text{MLP}(o_{i})))\right],$
where $\mathcal{M}$ is the set of masked interactions and $\sigma(q)=1/(1+e^{-q})$ is the sigmoid function. We apply the $L_{\text{pretrain}}$ in every sequence in $S_{\text{pretrain}}$.In our experiments, we test different mask ratios, which shows that the best mask ratio should be different for different datasets. In contrast to the mask item loss(Sun et al., 2019), when dealing with sequences that are annotated with multiple behaviors, it becomes relatively more feasible for the mask behavior loss to propagate and encompass these labeled behaviors.
4.3. Retriever
The retriever $R$ is responsible for retrieving similar context sequences for target interaction $d_{t}$. We use the feature part $x_{t}$ of $d_{t}$ as the query to utilize the search engine to find similar user contexts in the retrieval pool. This section will cover the detail of the datastore and search processes.
Datastore. The datastore serves as a database in which the keys correspond to all the interactions within the retrieval dataset denoted as $\mathcal{D}_{\text{retrieval}},$ while the values correspond to their associated context sequences. For the sake of expediency, we employ a pretrained encoder to perform offline encoding of all these sequences into their respective embedding representations while the key is still the sparse format for retrieval.
Figure 4. The datastore in LIFT. The primary keys within this datastore are intricately linked to the interactions in the dataset $\mathcal{D}_{\text{retrieval}}$, which represents the original data format. In parallel, the datastore houses values that align with the contextual information associated with these interactions. It is noteworthy that this contextual information is derived through an encoding process that applies a pre-trained encoder, denoted as $E$, to the context sequences.
The encoder $E_{\omega}$ encodes an interaction $d_{z}$’s context $c_{z}$’s history part $h_{z}$ and future part $f_{z}$ into two fixed-length vector representations respectively. Thus for any sample $d_{z}$ in $\mathcal{D}_{\text{retrieval}}$, we define the key-value pair $(k_{z},v_{z})$, where the key $k_{z}$ is the raw features $x_{z}$ of $d_{z}$ and the value $v_{z}$ is the encoded context embedding that contains a history sequence embedding $\mathbf{h}_{z}$ and a future sequence embedding $\mathbf{f}_{z}$. The datastore $(\mathcal{K},\mathcal{V})$ is the set of all key-value pairs constructed from all the samples in the retrieval dataset $\mathcal{D}_{\text{retrieval}}$:
(11) $(\mathcal{K},\mathcal{V})=\{(x_{z},(E_{\omega}(h_{z}),E_{\omega}(f_{z})))|d_{z}\in\mathcal{D}_{\text{retrieval}}\}~{}.$
Query. The retrieval process is to use the target sample’s raw features $x_{t}$ as the query to find the nearby keys in the datastore. We use the traditional BM25 (Robertson et al., 1995) algorithm to retrieve relevant keys, and return top $K$ ranked key-value pairs. In the BM25 algorithm, we treat any sample $x_{z}$ as the query and a key $x_{d}$ in the datastore as the document. As such, the ranking score can be calculated as
(12) $\text{RankScore}(x_{z},x_{d})$ $=\sum_{j=1}^{M}\text{IDF}(x_{j}^{z})\cdot\mathbf{1}(x_{j}^{z}=x_{j}^{d})~{},$
(13) $\text{IDF}(x_{j}^{z})$ $=\log\frac{N-N(x_{j}^{z})+0.5}{N(x_{j}^{z})+0.5}~{},$
where $\mathbf{1}(\cdot)$ is the indicator function, $N$ is the number of data samples in $\mathcal{D}_{\text{retrieval}}$, and $N(x_{j}^{z})$ is the number of data samples that have the categorical feature value $x_{j}^{z}$.
Let $\hat{\mathcal{K}}$ be the retrieved keys and $\hat{\mathcal{V}}$ be the retrieved values from the retrieved set. After the retrieval process, we obtain the top-$K$ relevance pairs in the datastore.
4.4. Predictor
After the retrieval process, we obtain the context of the target sample $x_{t}$. The contextual information includes the user $u_{t}$’s history and both the history and future information of the target sample’s relevant interactions. As shown in Figure 3 right part of stage 2, in the predictor, we use both the contextual information and the embedding of $x_{t}$ to produce the final label prediction $\hat{y_{t}}$.
Key-based Attention Aggregation. From the retriever, we get the context embedding set $\hat{\mathcal{V}}$ of relevant interactions. Before feeding to the final prediction layer, we use the key-based attention mechanism to aggregate the retrieved $\hat{\mathcal{K}}$ and $\hat{\mathcal{V}}$. Let $x_{i}$ be the $i$-th sample in the key set $\hat{\mathcal{K}}$. We use an embedding layer to convert $x_{i}$ into a $M\times w$-dimensional dense vector $\mathbf{x}_{i}$ and we use the same embedding layer to map $x_{t}$ into a $M\times w$-dimensional dense vector $\mathbf{x}_{t}$. The key attention weight is defined as
(14) $\alpha_{i}=\frac{e^{\mathbf{x}_{i}^{T}{W}\mathbf{x}_{t}}}{\sum_{j=1}^{K}e^{\mathbf{x}_{j}^{T}{W}\mathbf{x}_{t}}}~{},$
where ${W}\in\mathbb{R}^{Mw\times Mw}$ is the attention layer parameter matrix. The allocation of attention weights to value sets is determined based on their respective key attention weights. Thus we could use $\bf{\alpha}$ to aggregate the retrieved key set $\hat{\mathcal{K}}$ and value set $\hat{\mathcal{V}}$. For the retrieved samples in $\hat{\mathcal{K}}$, the aggregated vector $\mathbf{x}_{r}$ can be written as $\mathbf{x}_{r}=\sum_{i=1}^{K}\alpha_{i}\cdot\mathbf{x}_{i},~{}~{}x_{i}\in\hat{\mathcal{K}}$. For the history part of the retrieved context embeddings $\hat{\mathcal{V}}$, the aggregated embedding $\mathbf{e}_{r}^{h}$ can be written as $\mathbf{e}_{r}^{h}=\sum_{i=1}^{K}\alpha_{i}\cdot\mathbf{h}_{i},~{}~{}\mathbf{h}_{i}\in\hat{\mathcal{V}}$. Similarly, for the future part of the retrieved context embeddings $\hat{\mathcal{V}}$, the aggregated embedding $\mathbf{e}_{r}^{f}$ can be written as $\mathbf{e}_{r}^{f}=\sum_{i=1}^{K}\alpha_{i}\cdot\mathbf{f}_{i},~{}~{}\mathbf{f}_{i}\in\hat{\mathcal{V}}$.
Prediction. In the final prediction layer, we first use a factor interaction layer to perform the inter-sample interaction for the target sample and the retrieved key set. Since these two vectors are categorical data, one could use high-order interaction further explore the useful patterns, but this is out of the scope of this paper. Finally, we feed the embedding of the target sample, retrieved keys, and retrieved values to an MLP to make the label prediction.
After the embedding layer, each sample $x_{z}$ in $\mathcal{D}$ can be represented as a concatenation of $M$ feature embeddings:
(15) $\mathbf{x}_{z}=[{\mathbf{a}_{1}^{z}},\ldots,{\mathbf{a}_{M}^{z}}]~{},$
where $\mathbf{a}_{i}^{z}$ represents the $i$-the feature’s embedding of the sample $x_{z}$. We use an interaction layer to compute the combination of these features as
(16) ${\mathbf{x}_{\text{inter}}^{z}}=[\text{inter}(\mathbf{a}_{1}^{z},\mathbf{a}_{2}^{z}),\text{inter}(\mathbf{a}_{1}^{z},\mathbf{a}_{3}^{z}),\ldots,\text{inter}(\mathbf{a}_{M-1}^{z},\mathbf{a}_{M}^{z})]~{},$
where the $\text{inter}(\cdot)$ represents the interaction function. We could use different interaction functions such as inner product (Guo et al., 2017), kernel product (Qu et al., 2018), or micro-network (Qu et al., 2018). We use the same architecture of the interaction layer for the target sample $x_{t}$ and aggregated vector $\mathbf{x}_{r}$. After $\mathbf{x}_{r}$ going through the interaction layer, we get the embedding $\mathbf{x}_{\text{inter}}^{r}$. Finally, we concatenate all the vectors from the framework as
(17) $\text{hidden}=\text{concat}[\mathbf{x}_{\text{inter}}^{t},\mathbf{x}_{\text{inter}}^{r},\mathbf{e}_{r}^{f},\mathbf{e}_{r}^{h},\mathbf{h}_{t}]~{},$
where $\mathbf{h}_{t}$ is the user $u_{t}$’s history representation. As such, the label final predicted $\hat{y}_{z}$ is written as
(18) $\hat{y}_{z}=\sigma(\text{MLP}(\text{hidden}))~{}.$
Given the prediction is a binary classification for CTR prediction, we use the binary cross-entropy loss between the label $y_{t}$ and prediction $\hat{y}_{t}$ for training.
4.5. Time complexity & Speedup
In this section, we provide a brief time complexity analysis of LIFT framework and then discuss the speedup feasibility.
Time complexity. We aim to provide a comprehensive analysis of the time complexity within our proposed framework, which consists of two major components.
The first component addresses the pretraining phase targeted at context sequences. In this phase, the dataset $D_{\text{retrieval}}$ is segmented into a set of sequences, each of which is characterized by a length $L$. Both training and inference stages involve operations on each sequence with a time complexity of $O(1)$. Consequently, the overall time complexity during the pretraining phase is $O(|D_{\text{retrieval}}|/L)$.
The second component pertains to the complexity of the main framework’s algorithm, which diverges from traditional neural network methodologies by incorporating a retrieval process during both training and inference phases. We commence by examining the complexity of this retrieval process. Here, $|D_{\text{retrieval}}|$ represents the quantity of samples, and $U$ denotes the total number of unique features encountered in $D_{\text{retrieval}}$. Note that the mean length of the posting lists in the inverted index is $|D_{\text{retrieval}}|/U$. As explained in Section 4.3, the retrieval operation, which encompasses retrieving all posting lists of features in $x_{t}$, necessitates a time complexity of $O(F)$, where $F$ symbolizes the number of features in the query. This phase is deemed a constant time operation. The average count of samples retrieved is $F\cdot|D_{\text{retrieval}}|/U$, and the complexity of the ranking operation scales linearly with the number of retrieved samples, indicated as $O(F\cdot|D_{\text{retrieval}}|/U)$. Therefore, the total time complexity of the retrieval process is $O(F)+O(F\cdot|D_{\text{retrieval}}|/U)=O(F\cdot|D_{\text{retrieval}}|/U)$. Besides the retrieval process, the neural network computation time complexity of the predictor is about the same with the other models such as RIM (Qin et al., 2021) or DIN (Zhou et al., 2018b). The detailed architecture acceleration would be an effective factor for the model inference effieiency.
Speedup. To enhance the efficiency of the retrieval process, we may adopt a method of storing the retrieval outcomes, eliminating the need to conduct the retrieval procedure during online inference. This approach shifts the retrieval process to an offline setting. For instance, in recommender systems, this method involves storing the retrieval results for all (or frequent) users, along with their respective recalled candidate items. This method represents a balance between the complexities of space and time, optimizing resource utilization by trading off storage space against processing speed. On the other hand, a recent work (Zheng et al., 2024) has proposed the use of knowledge distillation to distill the datastore of a recommender system into a neural network module, thereby fully avoiding the time and disk space overhead brought by the retrieval module. Thus, it is highly potential that the efficiency of the retrieval process will be sufficiently optimized to make LIFT practically deployable.
5. Experiments
This section starts with five research questions (RQs), which we use to guide the experiments and discussions. [^1] We provide the experiment code with running instructions on Anonymous GitHub (https://anonymous.4open.science/r/LIFT-277C/).
  • RQ1
    Does LIFT achieve the best performance?
  • RQ2
    Does future data benefit the final result?
  • RQ3
    How does the proposed pretraining method have a positive impact on the prediction performance?
  • RQ4
    How do the hyperparameters ($L$ and $K$) influence the performance of LIFT?
  • RQ5
    Do different encoders and mask label ratios have different impacts on context representation learning?
  • RQ6
    How is the training and inference efficiency of LIFT and whether is it potential to be deployed online?
5.1. Datasets
We evaluate the performance of LIFT by conducting experiments for CTR prediction tasks on three large-scale real-world datasets, i.e., Taobao, Tmall, and Alipay [^2] https://tianchi.aliyun.com/dataset/x, where ‘x’ is ‘649’, ‘42’, and ‘53’ for Taobao, Tmall, and Alipay, respectively.. We count the number of instances, fields, and users of these datasets, which is shown in Table 2.
Datasets # Instances # Fields # Users
Taobao 100,150,807 4 987,994
Tmall 54,925,331 9 424,170
Alipay 35,179,371 6 498,308
Table 2. Dataset statistics.
Figure 5. Dataset split scheme.
As shown in Figure 5, following the experiment setup of (Qin et al., 2021), we split each dataset into three parts, i.e., retrieval set, train set, and test set, based on the order of the global timestamps. Specifically, the retrieval set consists of the earliest data instances, the test set comprises the latest data instances, and the remaining intermediate data instances are allocated to the train set. As for the pretraining, we use the retrieval set to pretrain the encoder. For the non-retrieval baseline models, we merge the retrieval set and the training set as the final training set.
5.2. Evaluation Metrics
We choose the widely used metrics area under the ROC curve (AUC) and negative log-likelihood (LogLoss) to evaluate the performance. We also conduct a significance test on the metrics between LIFT and the best performed baseline, marked with “*” for positive results.
Models Taobao Tmall Alipay
LogLoss AUC Rel. Impr. LogLoss AUC Rel. Impr. LogLoss AUC Rel. Impr.
GBDT 0.6797 0.6134 44.39% 0.5103 0.8319 11.13% 0.9062 0.6747 30.49%
DeepFM 0.6497 0.6710 32.00% 0.4695 0.8581 7.74% 0.6271 0.6971 26.29%
FATE 0.6497 0.6762 30.98% 0.4737 0.8553 8.09% 0.6199 0.7356 19.68%
BERT4Rec 0.6356 0.6852 29.26% 0.4017 0.8981 2.94% 0.6024 0.7321 20.26%
DIN 0.6086 0.7433 19.16% 0.4292 0.8796 5.10% 0.6044 0.7647 15.13%
DIEN 0.6084 0.7506 18.00% 0.4445 0.8838 4.61% 0.6454 0.7502 17.36%
SIM 0.5795 0.7825 13.19% 0.4520 0.8857 4.38% 0.6089 0.7600 15.84%
UBR 0.5432 0.8169 8.42% 0.4368 0.8975 3.01% 0.5747 0.7952 10.71%
RIM 0.4644 0.8563 3.43% 0.3804 0.9138 1.17% 0.5615 0.8006 9.97%
DERT 0.4486 0.8647 2.42% 0.3585 0.9200 0.4% 0.5319 0.8087 8.86%
LIFT w/o pretrain 0.4369* 0.8727* 1.49% 0.3509* 0.9236* 0.10% 0.4707* 0.8572* 2.71%
LIFT 0.4129* 0.8857* - 0.3489* 0.9245* - 0.4361* 0.8804* -
Table 3. Performance comparison of CTR prediction task baselines. GBDT and DeepFM are the traditional methods. Others are sequential modeling methods. For a fair comparison, traditional models are trained on both the retrieval set and the training set. The best results are in bold fonts while the second best results are underlined. “Rel. Impr.” of each row means the relative AUC improvement of LIFT against the baseline. Improvements are statistically significant with $p<0.01$.
Encoder RNN RNN(pretrain) Transformer Encoder Transformer Encoder (p) Transformer Decoder Transformer Decoder (p)
LogLoss AUC LogLoss AUC LogLoss AUC LogLoss AUC LogLoss AUC LogLoss AUC
Taobao 0.4434 0.8667 0.4360 0.8715 0.4432 0.8684 0.4454 0.8679 0.4369 0.8727 0.4129 0.8857
Tmall 0.3730 0.9132 0.3637 0.9175 0.3539 0.9223 0.3528 0.9227 0.3509 0.9236 0.3489 0.9245
Alipay 0.4937 0.8391 0.4856 0.8468 0.4807 0.8504 0.4707 0.8572 0.4730 0.8556 0.4361 0.8804
Table 4. The influence of different encoders of LIFT on prediction performance. “(p)” means the encoder is pretrained.
Dataset No Context Future Only History Only Future & History
LogLoss AUC LogLoss AUC LogLoss AUC LogLoss AUC
Taobao .4644 .8563 .4417 .8682 .4405 .8694 .4129 .8857
Tmall .3804 .9138 .3606 .9191 .3588 .9199 .3489 .9245
Alipay .5615 .8006 .4781 .8525 .4780 .8527 .4361 .8804
Table 5. The influence of the usage of future data in LIFT.
5.3. Compared Methods
We compare LIFT with nine solid baselines that can be categorized into three groups. The first group is traditional tabular models, which do not utilize the sequential or retrieval mechanism. This group includes GBDT (Chen et al., 1996), a widely-used gradient-boosted trees model, and DeepFM (Guo et al., 2017), a factorization-machine based deep neural network. The second group consists of end-to-end sequential deep models, such as DIN (Zhou et al., 2018b) and DIEN (Zhou et al., 2019) that are attention-based recurrent neural networks for CTR prediction, and BERT4Rec (Sun et al., 2019) that is a Transformers-based sequential recommendation model. The third group includes retrieval-based models, such as SIM (Qi et al., 2020), UBR (Qin et al., 2020), RIM (Qin et al., 2021) and DERT(Zheng et al., 2023). Additionally, FATE (Wu et al., 2021) is a tabular data representation learning model that can be viewed as a random retrieval method for prediction.
Additionally, FATE (Wu et al., 2021) is a tabular data representation learning model, which accepts a random mini-batch of data instances as the entire input and predicts all its elements simultaneously, enabling interactions between the samples in the mini-batch. This approach can be viewed as leveraging a random retrieval method for prediction.
5.4. Overall Performance (RQ1)
The overall performance comparison result is provided in Table 3, where we have the following observations: (i) LIFT consistently outperforms all the nine baselines. Specifically, compared to the best baseline RIM, LIFT achieves the relatively improved AUC by 3.43%, 1.17% and 9.97% on Taobao, Tmall, and Alipay, respectively. This clearly demonstrates the effectiveness of context information learned and retrieved in LIFT framework. (ii) Without the pretraining, LIFT still consistently outperforms the best baseline, which indicates that the raw information of future and history still yields a significant improvement on the prediction performance. (iii) The retrieval-based methods LIFT and RIM are superior to the traditional methods and sequential CTR models, which means that the retrieval methods are able to make better use of the context information.
5.5. Further Analysis
We further study the effectiveness of important modules in LIFT, i.e., the usage of future context, the pertaining, the encoder architecture, and the hyperparameters.
Future Context (RQ2). Table 5 compares the different parts of the context information’s impact on the final prediction performance. We can observe that both the history information and future information have a large impact on the final prediction. If we only use part of the context information, such as future only or history only, we will get a worse result than using both of them, which indicates that the future part of the context provides different information compared with the history part. This observation is totally neglected by previous works.
Pretraining (RQ3). From the last two rows of Table 3, we can find that over the three datasets, the model with pretraining performs better than those not, which clearly demonstrates the efficiency of pretraining. The enhanced performance of the pretrained encoder implies that relying solely on the label provided by the target sample is insufficient for effectively capturing contextual information. To improve the modeling of context, it is imperative to harness the inherent information embedded within the context itself.
Hyperparameter in the Retriever (RQ4). In the LIFT framework, $L$ means the context sequence length and $K$ means the retrieved samples count. From Figure 7 and Figure 8, we can find that with the increase of $K$, the final prediction result first goes up and then down. The curve indicates that along with with the $K$, the information first increases, and then more noises are introduced to become the dominant part. The AUC curves of $L$ show the similar trends of $K$, i.e., as $L$ increases, the AUC first gets better and then drops down. We think it is the same reason as for $K$. Moreover, because of the limitation of the GPU resources, we only conduct this hyperparameter study on Taobao and Alipay. and we could only increase the $L$ to 70 in the history length of Taobao, where the downtrend just started.
Figure 6. Performance w.r.t. different pretraining mask rates.
Figure 7. Hyperparameters study of LIFT on Taobao.
Figure 8. Hyperparameters study of LIFT on Alipay.
The Choice of the Encoder (RQ5). The prediction performance of LIFT with different encoders is provided in Table 4, where we can observe that the Transformer decoder based encoder yields the best performance on both non-pretraining and pertraining settings. The reason could be that the Transformer decoder could be aware of the position of the sequence better. This justifies our choice of architecture implementation of the encoder. Further, from Figure 6, we observe the mask ratio’s influence on the prediction performance. Usually, the traditional models that encode the user interaction sequence only use the last label as the supervised signal. From the result, we can find that in all three datasets, such $1$ label supervised strategy is not the best choice. The best choices mask label rate is different over datasets where the number is usually larger than 50%.
Time Efficiency (RQ6). To evaluate the time efficiency of LIFT, we compare the real inference time of it against other three mainstream models in different types. As shown in Figure 9, LIFT does not introduce significant overload compared with the basic retrieval-based model RIM. With offline pertaining, LIFT acts far more efficiently than the transformer-based model BERT4Rec. It is potential to deploy LIFT to the real recommender systems if some speedup schemes are adopted, such as offline retrieval and distillation.
Figure 9. Inference time on Alipay.
Training behavior. The pretraining and training curves of LIFT on three datasets are presented in Figure 10 and 11. In the LIFT model, there are two modules that need to be trained. One is the encoder module used to encode the user sequence. The other is the predictor module that makes prediction by combining user history and information from similar historical and future contexts.
In the LIFT model, our training is divided into two steps: first, we train the encoder module, followed by the predictor module. In Figure 10, we can see that the pre-training of the encoder module converges smoothly across all three datasets. In Figure 11, we observe that the inclusion of the retrieval module makes the neural network’s training curve appear steeper, but it ultimately converges. The curves demonstrate that the pretraining and training processes of LIFT are stable and effective.
Figure 10. Pretraining curves of LIFT.
Figure 11. Training curves of LIFT.
6. Conclusion
In this work, we propose a retrieval-based framework called LIFT to better utilize the context information of the current user interaction. We are the first work to include the future information as the part of the context without data leakage. Moreover, we use a pretraining method to mine the information in the context sequences better and propose a novel mask behavior loss. The performance of the LIFT framework shows that both history and future information yield a significant improvement in the CTR prediction performance. Also, from the comparison of the pretraining method, we find that the context pretraining is a potential way to further improve the prediction performance of a variety of sequential recommendation models. In the future, we will investigate deeper on context representation learning with more sophisticated retrieval methods and the speedup schemes to deploy LIFT to real-world recommenders.